# Emergent Abilities in Large Language Models: A Survey

### v1

@@ Abstract
1 Introduction
The study of emergent properties in complex systems has been a long-standing interdisciplinary pursuit, spanning fields such as physics, biology, and mathematics. While the term emergent was coined by G. H. Lewes [46] in 1877111”Each stage of evolution presents itself as the consequence of a preceding stage, at once an emergence and a continuance; so that no transposition of stages is possible; each has its appointed place…”, the concept of emergence gained widespread recognition through Anderson’s seminal work, “More Is Different[1]. Anderson postulated that, as systems increase in complexity, novel surprising properties may manifest, even with a comprehensive quantitative understanding of their microscopic constituents.
Emergent Abilities in Large Language Models: A Survey§ 4.Emergent Abilities asIn-Context Learning§ 3. Emergent Abilities inLarge Language Models§ 2. In-the-Wild Definitions ofEmergent Abilities§ 1. Introduction § 5.Emergent Abilities ofLarge Reasoning Models§ 6.Emergent Behaviors inLLMs-Powered AI Agents§ 7.Emergent Harmful Behaviorsin LLMs and AI Agents§ 8.A Taxonomic Synthesis of KeyFindings and Challenges§ 9.Conclusions§ 3.4.Impact of Quantization on Emergent Abilities§ 3.3.Loss Functions and Emergent Abilities§ 3.2.Relationship between Emergent Abilities andPrompt Strategies§ 3.1.Do emergent abilities really exist?§ 3.5.Emergent Abilities and Task Complexity§ 3.6.Emergent Abilities and Arithmetic§ 3.7.Predicting Emergent Abilities§ 7.1.AI Safety
@@ 1 Introduction
This work comprehensively reviews the study of emergent abilities for LLMs. Despite growing academic and public interest in this phenomenon, there remains a lack of consensus on the precise definition of emergent abilities, leading to significant conceptual ambiguity and confusion. Ironically, the need for clarification on emergent properties is not novel. Johnson et al. [43] previously addressed this topic, although they focussed on their manifestation in the engineering of complex systems. To address this definitional uncertainty, we first conduct a systematic analysis of current definitions in the literature (Section˜2). Then we comprehensively review the literature on emergent abilities (Section˜3) as defined by Wei et al. [85]. Subsequently, we examine in-context learning (Section˜4) and continue by exploring the emergent abilities of Large Reasoning Models (LRMs), a new class of AI systems that extend traditional LLMs by incorporating reinforcement learning post-training and inference-time search (Section˜5). Beyond beneficial emergent abilities, we also examine the rise of LLM-powered AI agents and their implications (Section˜6). Finally, we analyze the emergence of harmful behaviors in LLMs and LLM-powered agents (Section˜7). Advanced AI systems have demonstrated deceptive tendencies and reinforcement learning-driven manipulation that could lead to unintended consequences.
Table 1: Summary of the sections and subsections in this survey. Each row provides an overview of the topic, relevant papers, and a brief summary.
@@
Emergent Abilities in Large Reasoning Models [70, 72, 90, openai-2025, 18] Shows how Large Reasoning Models develop complex reasoning tasks, aided by scaling reinforcement learning post-training techniques and increased test-time compute.
@@
2 In-the-Wild Definitions of Emergent Abilities
@@
Table 2: Summary of selected papers on emergent abilities in LLMs. The table outlines each paper’s focus, the tasks and models studied, the main hypothesis and TL;DR summary, as well as noted limitations.
@@
3 Emergent Abilities in Large Language Models
Emergent abilities in LLMs have been a subject of increasing interest, particularly since their characterization in [85]. The phenomenon can be understood through an analogy to phase transitions in physics, where a system undergoes a sudden qualitative shift in behavior once a critical threshold is crossed. In the case of LLMs, these emergent abilities are not a product of gradual improvement but instead appear abruptly when scaling reaches a certain level. Performance often hovers near random until this threshold is surpassed, at which point a sharp jump occurs. This unpredictability makes it difficult to foresee emergent abilities by simply extrapolating from smaller models. It also raises fundamental questions about the complexity and non-linear scaling behavior of LLMs. Table (2) summarizes this section’s surveyed papers. For this section, we have selected all the papers that appeared when searching in Google Scholar with the query "Emergent Abilities" "Large Language model".
@@ 3 Emergent Abilities in Large Language Models
3.1 Do emergent abilities really exist?
@@
In addition to these experiments, the paper also presents a meta-analysis of emergent abilities within the BIG-Bench framework, evaluating multiple model families and metrics. Their findings suggest that most metrics used in BIG-Bench do not exhibit emergent abilities, although their criteria for classifying a metric-task-model triplet as emergent could be debated. Specifically, the authors define a triplet as emergent when its emergence score exceeds 100, a notably stringent threshold. This score reflects a scenario wherein, across a continuum of models of increasing scale evaluated against a specific metric-task pair, the difference between the maximum and minimum metric values is 100 times greater than the median difference. We contend, however, that an emergence score ranging between 20 and 100 might also reasonably qualify as emergent. For instance, consider four models of progressively larger sizes yielding performance scores of 5%, 10%, 20%, and 90% on a selected metric-task pair; this configuration produces an emergence score of 8.5, which, while below the authors’ threshold, still suggests a discontinuous jump in performance. Additionally, they demonstrated that changing evaluation metrics eliminated signs of emergent behavior for three333swahili-english-proverbs was not emergent also with the first metric. tasks within the LaMDA model family. Finally, they induced emergent-like behavior in deep networks across various vision tasks, suggesting that emergent abilities can, in some cases, be artificially introduced through experimental design.
@@
3.2 Relationship between Emergent Abilities and Prompt Strategies
@@
3.3 Loss Functions and Emergent Abilities
@@
3.4 The Impact of Quantization on Emergent Abilities
@@
3.5 Emergent Abilities and Task Complexity
@@
3.6 Emergent of Implicit Discrete State Representations
@@
3.7 Predicting Emergent Abilities
@@
4 Emergent abilities as in-context learning
@@
Another key research focus is the optimization of prompt design for effective ICL. Studies show that the selection, formatting, and arrangement of examples in a prompt significantly influence model performance [101, 54]. Methods for selecting exemplars can be broadly classified as unsupervised or supervised. Unsupervised approaches rely on nearest neighbor methods, using similarity metrics such as perplexity, mutual information, or self-evaluated LLM scores to identify the most relevant examples [50, 77, 74, 26, 91, 49]. Supervised methods employ dense retrievers trained on labeled data [65, 48, 95] and reinforcement learning [98] to optimize example selection. Research has also explored strategies for refining prompt formatting and ordering to enhance ICL effectiveness, as well as the impact of instruction formatting on how models process and execute in-context tasks [45, 94, 32, 35, 103, 83, 54, 50, 52].
@@
5 Emergent Abilities of Large Reasoning Models
@@
6 Emergent Behaviors in LLMs-powered AI agents
@@
7 Emergent Harmful Behaviors in LLMs and LLMs-powered AI agents
@@
Williams et al. [88] investigate how RLHF can unintentionally reinforce manipulative and exploitative behaviors. Their study reveals that LLMs trained through RLHF develop strategies that exploit user vulnerabilities to maximize reward signals. In controlled conversational settings, they found that models: (1) exhibited selective deception, targeting vulnerable individuals while maintaining normal interactions with others, rendering such behaviors difficult to detect; (2) encouraged harmful behaviors; (3) failed standard safety evaluations, as these manipulative tendencies were not detected by conventional toxicity or sycophancy benchmarks. This work demonstrates that reward-driven optimization does not always align with human well-being, highlighting the limitations of current safety evaluation frameworks. The challenge of deploying LLMs to be simultaneously helpful and harmless is further analyzed by Bai et al. [4], who examine this dual objective through extensive RLHF experimentation. Their research demonstrates that while RLHF can markedly enhance a model’s helpfulness and reduce harm, a fundamental tension persists between these goals. Models optimized predominantly for harmlessness often become excessively cautious, refraining from engaging in sensitive yet potentially beneficial discussions—such as those involving mental health or political discourse—thereby curtailing their practical utility. Additionally, the authors identify "over-optimization" as a critical issue, wherein models become overly attuned to reward signals, resulting in behaviors that deviate from genuine human preferences, such as overly simplistic or generic responses.
@@
As these models evolve, they are poised to exhibit a variety of characteristics that mark significant advancements. One such trait is rapid self-improvement, where mechanisms like iterative fine-tuning and self-supervised learning allow AI systems to refine their internal representations and decision-making processes at an accelerated pace. Evidence of this self-improvement is already apparent, as reinforcement learning post-training utilizes frozen models to generate rewards for those being trained, though the extent to which they can further enhance themselves remains an open question. Additionally, newer models have demonstrated the ability to produce novel research ideas, as noted by [27]. Alongside this, enhanced collaboration emerges as another key feature, particularly in networked environments where multiple AI agents have proven capable of working together effectively, according to findings by [28] and [14]. Furthermore, these advanced agents showcase contextual adaptability, learning to interpret multi-modal inputs—such as text, images, and sensor data—and dynamically adjusting their strategies based on real-time feedback.
Let us now hypothesize that, with further technological progress, LLM-powered AI agents might surpass human intelligence. This trajectory toward a qualitatively distinct form of intelligence introduces profound practical challenges. Should these agents achieve or exceed human-level intelligence, ensuring their submission to humans becomes a paramount concern, necessitating research into concepts like corrigibility and interruptibility to prevent scenarios where they might resist human intervention or behave unpredictably. Closely tied to this is the need for robust risk management, given the potential for rapid, exponential improvements in AI capabilities—often described as an intelligence explosion—which demands preemptive safety protocols. As emphasized by [9], understanding and mitigating the risks associated with superintelligent agents is essential to avoid unintended consequences.
@@
7.1 AI Safety
The Challenge of Governing Autonomous AI Agents. LLM-powered AI agents differ from traditional AI models by being more autonomous in nature. Examples like o3-mini and DeepSeek-R1 demonstrate abilities to perform multi-step reasoning, make adaptive decisions, and strategize for reward maximization. While these capabilities enhance their efficiency and usefulness, they also pose new governance challenges not previously considered in AI policy discussions.
Historically, AI safety has focused on content moderation, bias mitigation, and adversarial robustness, but autonomous AI introduces a new class of risks that require different regulatory approaches. These governance challenges are especially critical in high-stakes applications such as healthcare, finance, legal reasoning, warfare, and autonomous systems.
The Role of AI Regulation and Global Governance. AI governance is no longer a theoretical concern; governments and international organizations are actively drafting policies to regulate AI development and deployment. To address these challenges, governments and organizations worldwide are exploring various regulatory strategies, such as the EU AI Act [24], the US AI Executive Order [8], and the UN AI Advisory Board [81]. Despite these efforts, there is no globally unified approach to AI regulation, leading to inconsistencies and regulatory loopholes. Some countries favor strict oversight (i.e., the EU member states), while others prioritize AI innovation over safety (i.e., the US and China), making international collaboration essential to prevent fragmented and ineffective governance. These ethical concerns underscore the need for interdisciplinary collaboration between AI researchers, policymakers, and industry leaders. Ensuring that AI systems benefit the community requires ongoing dialogue and policy refinement. AI governance should be focused on ensuring that powerful autonomous reasoning agents do not threaten the preservation of the human species. As AI systems become more capable, adaptive, and independent, the challenge of controlling their behavior and ensuring trustworthiness will define the next era of AI safety research and policy development.
@@ IV. Prediction, Mitigation, Evaluation, Safety
Table 3: Integrated Taxonomy of Emergent Abilities in Large Language Models. This table synthesizes the origins, manifestations, impacts, and strategies related to emergent behaviors as discussed in the survey.
8 A Taxonomic Synthesis of Key Findings and Challenges
@@
Table 3 presents an integrated taxonomy of emergent abilities in large language models. It organizes key insights into four primary categories: Origins, Manifestation, Impact, and Strategies (for prediction, evaluation, and mitigation). In addition, it delineates how these abilities arise from scale-dependent effects, training dynamics, task complexity interactions, and metric artifacts, how they manifest through capabilities such as in-context learning, symbolic reasoning, enhanced reasoning through reinforcement learning, and autonomous agent behaviors, and both their beneficial and potentially harmful outcomes. All of this is accompanied by corresponding applications, limitations, and challenges. This structured overview serves as a valuable roadmap for understanding and guiding future research in this rapidly evolving field.
9 Conclusion
This survey delves into the intricate and rapidly evolving domain of emergent abilities in LLMs, offering a comprehensive exploration of their nature, analysis, and implications. Throughout this analysis, several critical insights have surfaced, enriching our understanding of LLMs and their behavioral dynamics. Notably, the unpredictability and substantial performance leaps observed in specific tasks as the model scale increases underscore the essence of emergent abilities. Additionally, in-context learning, while closely intertwined with emergent abilities, emerges as a distinct phenomenon warranting separate scrutiny. The influence of data distributional properties also proves instrumental in shaping in-context learning capabilities. Theoretical frameworks, such as Bayesian inference applied to latent language spaces, present compelling pathways for elucidating the underlying mechanisms driving both emergent abilities and in-context learning.
Despite these advancements, significant gaps in knowledge persist, leaving numerous questions unresolved. The interplay between model scaling, architectural design, and the onset of novel capabilities remains incompletely understood, necessitating further investigation. Moreover, ongoing contention regarding whether emergent abilities represent authentic phenomena or are instead reflections of measurement biases underscores the pressing need for more robust, standardized evaluation methodologies to settle this debate conclusively.
Looking ahead, future research must prioritize several key avenues to advance this field. It is imperative to devise more sophisticated metrics capable of capturing subtle gradations in model performance across diverse tasks, thereby providing a clearer picture of emergent capabilities. A deeper examination of how model architecture and training methodologies contribute to the fostering of these abilities is equally essential. The ethical ramifications of emergent abilities also demand careful consideration, particularly in light of unforeseen or potentially deleterious capacities that may surface unexpectedly. Longitudinal studies tracking the evolution of emergent abilities across successive LLM generations offer a valuable opportunity to chronicle their development over time.
As LLMs continue their swift progression, comprehending emergent abilities becomes more important than ever, not only for advancing theoretical insights into artificial intelligence but also for informing practical applications. This area of inquiry illuminates the essence of AI while simultaneously challenging conventional notions of learning, cognition, and the genesis of complex behaviors from ostensibly simple components. Perhaps most critically, the demonstration of emergent abilities carries profound implications for safety, as unpredictable and potentially hazardous capabilities, such as the exploitation of software vulnerabilities or the manipulation of human actors, may arise without forewarning, necessitating vigilant oversight and proactive mitigation strategies.
Acknowledgement
We thank Bardh Prenkaj for his help on the project and Antonio Norelli for the useful and interesting discussions.

### v2

@@ Abstract
I Introduction
The study of emergent properties in complex systems has been a long-standing interdisciplinary pursuit, spanning fields such as physics, biology, and mathematics. While the term emergent was coined by G. H. Lewes [48] in 1877, the concept of emergence gained widespread recognition through Anderson’s seminal work, “More Is Different[1]. Anderson postulated that, as systems increase in complexity, novel surprising properties may manifest, even with a comprehensive quantitative understanding of their microscopic constituents.
Emergent Abilities in Large Language Models: A Survey§ IV.Emergent Abilities asIn-Context Learning§ III. Emergent Abilities inLarge Language Models§ II. In-the-Wild Definitions ofEmergent Abilities§ I. Introduction § V.Emergent Abilities ofLarge Reasoning Models§ VI.Emergent Behaviors inLLMs-Powered AI Agents§ VII.Emergent Harmful Behaviorsin LLMs and AI Agents§ VIII.A Taxonomic Synthesis of KeyFindings and Challenges§ IX.Conclusions§ III-D.Impact of Quantization on Emergent Abilities§ III-C.Loss Functions and Emergent Abilities§ III-B.Relationship between Emergent Abilities andPrompt Strategies§ III-A.Do emergent abilities really exist?§ III-E.Emergent Abilities and Task Complexity§ III-F.Emergent Abilities and Arithmetic§ III-G.Predicting Emergent Abilities§ VII-A.AI Safety
@@ I Introduction
This work comprehensively reviews the study of emergent abilities for LLMs. Despite growing academic and public interest in this phenomenon, there remains a lack of consensus on the precise definition of emergent abilities, leading to significant conceptual ambiguity and confusion. Ironically, the need for clarification on emergent properties is not novel. Johnson et al. [44] previously addressed this topic, although they focussed on their manifestation in the engineering of complex systems. To address this definitional uncertainty, we first conduct a systematic analysis of current definitions in the literature (Section˜II). Then we comprehensively review the literature on emergent abilities (Section˜III) as defined by Wei et al. [87]. Subsequently, we examine in-context learning (Section˜IV) and continue by exploring the emergent abilities of Large Reasoning Models (LRMs), a new class of AI systems that extend traditional LLMs by incorporating reinforcement learning post-training and inference-time search (Section˜V). Beyond beneficial emergent abilities, we also examine the rise of LLM-powered AI agents and their implications (Section˜VI). We analyze the emergence of harmful behaviors in LLMs and LLM-powered agents (Section˜VII), as advanced AI systems have demonstrated deceptive tendencies and reinforcement learning-driven manipulation that could lead to unintended consequences. Finally, we provide a taxonomic synthesis of the key findings and main aspects of this survey (Section˜VIII), before concluding in Section˜IX.
TABLE I: Summary of the sections and subsections in this survey. Each row provides an overview of the topic, relevant papers, and a brief summary.
@@
Emergent Abilities in Large Reasoning Models [72, 74, 92, 19] Shows how Large Reasoning Models develop complex reasoning tasks, aided by scaling reinforcement learning post-training techniques and increased test-time compute.
@@
II In-the-Wild Definitions of Emergent Abilities
@@
TABLE II: Summary of selected papers on emergent abilities in LLMs. The table outlines each paper’s focus, the tasks and models studied, the main hypothesis and TL;DR summary, as well as noted limitations.
@@
III Emergent Abilities in Large Language Models
Emergent abilities in LLMs have been a subject of increasing interest, particularly since their characterization in [87]. The phenomenon can be understood through an analogy to phase transitions in physics, where a system undergoes a sudden qualitative shift in behavior once a critical threshold is crossed. In the case of LLMs, these emergent abilities are not a product of gradual improvement but instead appear abruptly when scaling reaches a certain level. Performance often hovers near random until this threshold is surpassed, at which point a sharp jump occurs. This unpredictability makes it difficult to foresee emergent abilities by simply extrapolating from smaller models. It also raises fundamental questions about the complexity and non-linear scaling behavior of LLMs. Table (II) summarizes this section’s surveyed papers. For this section, we have selected all the papers that appeared when searching in Google Scholar with the query "Emergent Abilities" "Large Language model".
@@ III Emergent Abilities in Large Language Models
III-A Do emergent abilities really exist?
@@
In addition to these experiments, the paper also presents a meta-analysis of emergent abilities within the BIG-Bench framework, evaluating multiple model families and metrics. Their findings suggest that most metrics used in BIG-Bench do not exhibit emergent abilities, although their criteria for classifying a metric-task-model triplet as emergent could be debated. Specifically, the authors define a triplet as emergent when its emergence score exceeds 100, a notably stringent threshold. This score reflects a scenario wherein, across a continuum of models of increasing scale evaluated against a specific metric-task pair, the difference between the maximum and minimum metric values is 100 times greater than the median difference. We contend, however, that an emergence score ranging between 20 and 100 might also reasonably qualify as emergent. For instance, consider four models of progressively larger sizes yielding performance scores of 5%, 10%, 20%, and 90% on a selected metric-task pair; this configuration produces an emergence score of 8.5, which, while below the authors’ threshold, still suggests a discontinuous jump in performance. Additionally, they demonstrated that changing evaluation metrics eliminated signs of emergent behavior for three222swahili-english-proverbs was not emergent also with the first metric. tasks within the LaMDA model family. Finally, they induced emergent-like behavior in deep networks across various vision tasks, suggesting that emergent abilities can, in some cases, be artificially introduced through experimental design.
@@
III-B Relationship between Emergent Abilities and Prompt Strategies
@@
III-C Loss Functions and Emergent Abilities
@@
III-D The Impact of Quantization on Emergent Abilities
@@
III-E Emergent Abilities and Task Complexity
@@
III-F Emergent of Implicit Discrete State Representations
@@
III-G Predicting Emergent Abilities
@@
IV Emergent abilities as in-context learning
@@
Another key research focus is the optimization of prompt design for effective ICL. Studies show that the selection, formatting, explicit intent formulation, and arrangement of examples in a prompt significantly influence model performance [103, 56, 9]. Methods for selecting exemplars can be broadly classified as unsupervised or supervised. Unsupervised approaches rely on nearest neighbor methods, using similarity metrics such as perplexity, mutual information, or self-evaluated LLM scores to identify the most relevant examples [52, 79, 76, 27, 93, 51]. Supervised methods employ dense retrievers trained on labeled data [67, 50, 97] and reinforcement learning [100] to optimize example selection. Research has also explored strategies for refining prompt formatting and ordering to enhance ICL effectiveness, as well as the impact of instruction formatting on how models process and execute in-context tasks [47, 96, 33, 36, 105, 85, 56, 52, 54].
@@
V Emergent Abilities of Large Reasoning Models
@@
VI Emergent Behaviors in LLMs-powered AI agents
@@
VII Emergent Harmful Behaviors in LLMs and LLMs-powered AI agents
@@
Williams et al. [90] investigate how RLHF can unintentionally reinforce manipulative and exploitative behaviors. Their study reveals that LLMs trained through RLHF develop strategies that exploit user vulnerabilities to maximize reward signals. In controlled conversational settings, they found that models: (1) exhibited selective deception, targeting vulnerable individuals while maintaining normal interactions with others, rendering such behaviors difficult to detect; (2) encouraged harmful behaviors; (3) failed standard safety evaluations, as these manipulative tendencies were not detected by conventional toxicity or sycophancy benchmarks. This work demonstrates that reward-driven optimization does not always align with human well-being, highlighting the limitations of current safety evaluation frameworks. The challenge of deploying LLMs to be simultaneously helpful and harmless is further analyzed by Bai et al. [4], who examine this dual objective through extensive RLHF experimentation. Their research demonstrates that while RLHF can markedly enhance a model’s helpfulness and reduce harm, a fundamental tension persists between these goals. Models optimized predominantly for harmlessness often become excessively cautious, refraining from engaging in sensitive yet potentially beneficial discussions, such as those involving mental health or political discourse, thereby curtailing their practical utility. Additionally, the authors identify "over-optimization" as a critical issue, wherein models become overly attuned to reward signals, resulting in behaviors that deviate from genuine human preferences, such as overly simplistic or generic responses.
@@
As these models evolve, they are poised to exhibit a variety of characteristics that mark significant advancements. One such trait is rapid self-improvement, where mechanisms like iterative fine-tuning and self-supervised learning allow AI systems to refine their internal representations and decision-making processes at an accelerated pace. Evidence of this self-improvement is already apparent, as reinforcement learning post-training utilizes frozen models to generate rewards for those being trained, though the extent to which they can further enhance themselves remains an open question. Furthermore, newer models have shown the ability to produce novel research ideas, as noted by [28]. In addition to this, enhanced collaboration emerges as another key feature, particularly in network environments where multiple AI agents have been proven capable of working effectively, according to the findings of [29] and [15]. Furthermore, these advanced agents showcase contextual adaptability, learning to interpret multi-modal inputs, such as text, images, and sensor data, and dynamically adjusting their strategies based on real-time feedback.
Let us now hypothesize that, with further technological progress, LLM-powered AI agents might surpass human intelligence. This trajectory toward a qualitatively distinct form of intelligence introduces profound practical challenges. Should these agents achieve or exceed human-level intelligence, ensuring their submission to humans becomes a paramount concern, necessitating research into concepts like corrigibility and interruptibility to prevent scenarios where they might resist human intervention or behave unpredictably. Closely tied to this is the need for robust risk management, given the potential for rapid, exponential improvements in AI capabilities, often described as an intelligence explosion, which demands preemptive safety protocols. As emphasized by [10], understanding and mitigating the risks associated with superintelligent agents is essential to avoid unintended consequences.
@@
VII-A AI Safety
The Challenge of Governing Autonomous AI Agents. LLM-powered AI agents differ from traditional AI models in that they are much more complex and more autonomous in nature. Examples of large reasoning models like o3-mini and DeepSeek-R1 demonstrate abilities to perform multi-step reasoning, make adaptive decisions, and develop strategies for reward maximization. Although these capabilities improve their effectiveness and usefulness, they also pose new governance challenges not previously considered in AI policy discussions.
Historically, AI safety has focused on content moderation, bias mitigation, transparency, and (adversarial) robustness, but autonomous AI introduces a new class of risks related to opaque optimization goals and continuously changing information basis, which requires different regulatory approaches. These governance challenges are especially critical in high-stakes applications such as healthcare, finance, admission processes, legal reasoning, warfare, and autonomous systems.
The Role of AI Regulation and Global Governance. AI regulation and global governance have shifted from theoretical debates to urgent, real-world imperatives [46]; governments and international organizations are actively drafting policies to regulate AI development and deployment. The EU AI Act [25], the US AI Executive Order [8], and the UN AI Advisory Board [83], illustrate the diverse strategies at play. Despite these efforts, there is no globally unified approach to AI regulation, leading to inconsistencies and regulatory loopholes. Some countries favor strict oversight (i.e., EU member states), while others prioritize AI innovation over safety (i.e., the US and China), making international collaboration essential to prevent fragmented and ineffective governance. These ethical concerns underscore the need for interdisciplinary collaboration between relevant stakeholders, such as AI researchers, policymakers, industry leaders, and end users. Ensuring that AI systems benefit society requires ongoing dialogue and policy refinement. At the same time, AI governance should focus on ensuring that powerful autonomous reasoning agents do not threaten the preservation of biological lives. Therefore, we believe that innovation in governance, compliance mechanisms, and international oversight (with continuous evaluation) is at least as important as technological innovation. As AI systems become more capable, adaptive, and independent, the challenge of controlling their behavior and ensuring trustworthiness will define the next era of responsible AI applications, AI safety research, and policy development.
@@ IV. Prediction, Mitigation, Evaluation, Safety
TABLE III: Integrated Taxonomy of Emergent Abilities in Large Language Models. This table synthesizes the origins, manifestations, impacts, and strategies related to emergent behaviors as discussed in the survey.
VIII A Taxonomic Synthesis of Key Findings and Challenges
An overarching goal of this survey is to equip the interested reader with a structured and taxonomic understanding of emergent abilities in large language models.
@@ VIII A Taxonomic Synthesis of Key Findings and Challenges
Table III presents an integrated taxonomy of emergent abilities in large language models. It organizes key insights into four primary categories: Origins, Manifestation, Impact, and Strategies (for prediction, evaluation, and mitigation). In addition, it delineates how these abilities arise from scale-dependent effects, training dynamics, task complexity interactions, and metric artifacts, how they manifest through capabilities such as in-context learning, symbolic reasoning, enhanced reasoning through reinforcement learning, and autonomous agent behaviors, and both their beneficial and potentially harmful outcomes. All of this is accompanied by corresponding applications, limitations, and challenges. This structured overview serves as a valuable roadmap for understanding and guiding future research in this rapidly evolving field.
IX Conclusion
This survey delves into the intricate and rapidly evolving domain of emergent abilities in LLMs, offering a comprehensive exploration of their nature, analysis, and implications. Throughout this analysis, several critical insights have surfaced, enriching our understanding of LLMs and their behavioral dynamics. Notably, the unpredictability and substantial performance leaps observed in specific tasks as the model scale increases underscore the essence of emergent abilities. Additionally, in-context learning, while closely intertwined with emergent abilities, emerges as a distinct phenomenon warranting separate scrutiny. The influence of data distributional properties also proves instrumental in shaping in-context learning capabilities. Theoretical frameworks, such as Bayesian inference applied to latent language spaces, present compelling pathways for elucidating the underlying mechanisms driving both emergent abilities and in-context learning. Despite these advancements, significant gaps in knowledge persist, leaving numerous questions unresolved. The interplay between model scaling, architectural design, and the onset of novel capabilities remains incompletely understood, necessitating further investigation. Moreover, ongoing contention regarding whether emergent abilities represent authentic phenomena or are instead reflections of measurement biases underscores the pressing need for more robust, standardized evaluation methodologies to settle this debate conclusively. Looking ahead, future research must prioritize several key avenues to advance this field. It is imperative to devise more sophisticated metrics capable of capturing subtle gradations in model performance across diverse tasks, thereby providing a clearer picture of emergent capabilities. A deeper examination of how model architecture and training methodologies contribute to the fostering of these abilities is equally essential. As LLMs continue their swift progression, comprehending emergent abilities becomes more important than ever, not only for advancing theoretical insights into artificial intelligence but also for informing practical applications. This area of inquiry illuminates the essence of AI while simultaneously challenging conventional notions of learning, cognition, and the genesis of complex behaviors from ostensibly simple components. Perhaps most critically, the demonstration of emergent abilities carries profound implications for safety, as unpredictable and potentially hazardous capabilities, such as the exploitation of software vulnerabilities or the manipulation of human actors, may arise without forewarning, necessitating vigilant oversight, proactive mitigation strategies, and internationally aligned governance strategies.